19a-3(a-6)=66 whats a??
Answer by UJIMUSHKU(3)

19a-3(a-6)=66
multiply the 3 x (a-6)
19a-3a-18=66
combine like terms: Add 19a and -3a. Add 18 to both sides.
16a = 84
isolate the variable. Divide both sides by 16.
a = 84/16
reduce to a = 21/4